Solve and check the following systems by substitution.
(7x – 2y = 1)
(4x + 5y = 2)

Answer by mananth(16946)

7x – 2y = 1
7x=2y+1
/7
x=(2y+1)/7
4x + 5y = 2
...
4*(2y+1)/7 +5y=2
multiply by 7
4(2y+1)+7*5y=14
8y+4+35y=14
43y=10
y =10/43
...
plug value of y in (1)
7x-2*10/43=1
7x=1+20/43
7x=63/43
/7
x= 9/43